OpenStudy (anonymous):

Find the values of k such that f(x)=x^2+kx+k+3 has no real roots.

OpenStudy (raden):

hint : D < 0 D = discriminant = b^2 - 4ac

OpenStudy (raden):

did you get it, @rawritsmeli ?

OpenStudy (anonymous):

Could the answer be (0, squareroot 12) -thats interval notation btw...

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

f(x)=x^2+kx+k+3 here, a = 1, b = k, and c = k+3 now, the value of discriminant is D = b^2 - 4ac = k^2 - 4(1)(k+3) = k^2 - 4k - 12

OpenStudy (raden):

D < 0, it means k^2 - 4k - 12 < 0 (k -6)(k + 2) < 0 the value of k which makes the zeroes is k = 6 or k = -2 graph the lines of interval for k and give the sign negative or positive values for k ++++(-2)--------(6)++++++ chose the negative area, because the problem is < 0, so the solution for k be -2<k<6

OpenStudy (raden):

does that make sense ?? @rawritsmeli

OpenStudy (anonymous):

Woah!! yes! ok so i totally didn't think that k+3=c

OpenStudy (raden):

ok, just remember that the general equation of function quadratic is y = ax^2 + bx + c

OpenStudy (raden):

or f(x) = ax^2 + bx + c